# Rate-reliability functions

### v2

@@ III Bounding the rate-reliability function
We divide this section into three parts. The first part (Subsection III-A) presents our lower bound on the rate-reliability function for deterministic identification, expressed in terms of the packing number of a transformed output probability set, where the radius is determined by the error exponents. This is followed by an analysis of this bound in the regime of small error exponents. Similarly, the second part (Subsection III-B) provides the upper bound on the rate-reliability function using a covering number, along with its analysis for small error exponents. Finally, in Subsection III-C we analyse the rate bounds when the error exponents are restricted to certain suitable subsets in the vicinity of 00, raising our lower bound and decreasing our upper bound in the regime of small errors. Analogous to the results of [CDBW:DI_classical], this illuminates the separate roles of the lower and upper Minkowski dimension.
@@ III-A Lower bound on the rate-reliability function (coding)
We give our lower bound in Theorem 4, and analyse the regime of small error exponents in Corollary 5. In all the following sections we will use the shorthand $c=1/36K(|\mathcal{Y}|)$.
@@
For any $10<t<$ 1, and $E(n)>0$, there exists a DI code on block length $n$ with error exponents $E_{1}(n)\geq E(n)-\frac{1}{n}$ and $E_{2}(n)\geq E(n)-\frac{3}{n}$, and its rate lower-bounded
$\begin{split}R(n)&\geq(1-t)\log\left[\Pi_{\sqrt[4]{\frac{6E(n)}{ct^{2}}}}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\right]\\ &\phantom{\geq.}-H(t,1-t)-O\left(\frac{\log n}{n}\right).\end{split}$ (9)
@@
Inspired by the code construction in [CDBW:DI_classical], we use the typical sets of input strings as decoding elements, for which we know that Lemmas 2 and 3 give us upper bounds on the errors of first and second kind. We want to impose exponentially decreasing errors, construct a code, and then analyze the resulting rate. Looking at Lemma 2 it is immediate to see that for the error of first kind, the exponential decrease follows by letting $\delta:=\tau\sqrt{n}$, so that
$\lambda_{1}\leq 2\exp(-c\tau^{2}n).$ (10)
To bound the error of second kind, additional effort is required. We want to use Lemma 3, which requires a minimum separation $1-\frac{1}{2}\|W_{u_{j}}-W_{u_{k}}\|_{1}\leq\epsilon$ between any two different code words $u_{j}\neq u_{k}\in\mathcal{X}^{n}$. That is, the output probability distributions generated by code words need to form a packing with pairwise total variation distance being almost exponentially close to $1$. Following [CDBW:DI_classical], starting from
$1-\frac{1}{2}\|W_{x^{n}}-W_{{x^{\prime}}^{n}}\|_{1}\leq F\left(W_{x^{n}},W_{{x^{\prime}}^{n}}\right)=\prod_{i=1}^{n}F\left(W_{x_{i}},W_{x_{i}^{\prime}}\right),$ (11)
we study the negative logarithm of the left hand side. Letting $u_{j}=x^{n}=x_{1}\dots x_{n}$ and $u_{k}=x^{\prime n}=x^{\prime}_{1}\dots x^{\prime}_{n}$, we have
@@
Notice the important simplification that this achieves: before, we had to construct and analyze a packing of full sequences of length $n$ in total variation distance, and now we just need a letter-wise Euclidean packing to ensure a distance between output probability distributions exponentially close to 1. Let $\mathcal{X}_{0}$ be a maximum-size Euclidean packing in $\sqrt{\!\widetilde{\mathcal{X}}}$ of radius $\beta$ and $\Pi_{\beta}:=\Pi_{\beta}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)=|\mathcal{X}_{0}|$ the corresponding packing number.
Next we choose a maximum-size code $\mathcal{C}_{t}\subset\mathcal{X}_{0}^{n}$ such that the letters are elements of the packing and there is a minimum Hamming distance $d_{H}(x^{n},x^{\prime n})>tn$ between code words $x^{n}\neq x^{\prime n}\in\mathcal{C}_{t}$. Then, in the last sum in Eq. (12) at least $tn$ terms are $\geq 4\beta^{2}$ (and could be 00 in the others), therefore
@@
$|\mathcal{C}_{t}|\geq\Pi_{\beta}^{n(1-t)}2^{-nH(t,1-t)}.$
The Gilbert-Varshamov bound [Gilbert:combinatorics, Var:combinatorics] provides a similar (asymptotically equivalent) result, by constructing a linear code over the prime field $\mathbb{F}_{p}$, after reducing $\mathcal{X}_{0}$ to the nearest smaller prime cardinality $p\geq\frac{1}{2}|\mathcal{X}_{0}|$ (Bertrand’s postulate). It only remains to bound the entropy difference $H(W_{x^{n}})-H(W_{x^{\prime n}})$ to apply Lemma 3. As these entropies are in the interval $[0;n\log|\mathcal{Y}|]$, we can just partition the code into $S=\lceil n\log|\mathcal{Y}|\rceil$ parts $\mathcal{C}_{t}^{(s)}$ ($s=1,\ldots,S$) in such a way that all $j\in\mathcal{C}_{t}^{(s)}$ have $H(W_{u_{j}})\in[s-1;s]$, meaning that any two $j,k\in\mathcal{C}_{s}$ satisfy $|H(W_{u_{j}})-H(W_{u_{k}})|\leq 1$. We can now define $\mathcal{C}$ as the largest $\mathcal{C}_{t}^{(s)}$, which by the pigeonhole principle has $|\mathcal{C}|\geq|\mathcal{C}_{t}|/\lceil n\log|\mathcal{Y}|\rceil$ elements. We thus get the rate of the code as
@@
$\begin{split}\lambda_{2}&\leq 2\exp(-c\tau^{2}n)+3\exp(2\tau n-t\beta^{2}n)\\ &\leq 5\exp(-c\tau^{2}n),\end{split}$
where the last inequality is true if $-c\tau^{2}\geq 2\tau-t\beta^{2}$. The two error exponents in the present construction are: $E_{1}(n)\geq c\tau^{2}-\frac{1}{n}$ and $E_{2}(n)\geq c\tau^{2}-\frac{3}{n}$. Solving the quadratic constraint we obtain
$\tau\leq\frac{-2+\sqrt{4+4ct\beta^{2}}}{2c}=\frac{\sqrt{1+ct\beta^{2}}-1}{c}.$
We may assume w.l.o.g. $ct\beta^{2}\leq 1$ (see the definition of $c$, the range of $t$ and the meaningful range of $\beta$; otherwise our rate lower bound is trivial), hence by the concavity of the square root the above condition will hold for $\tau=\left(\sqrt{2}-1\right)t\beta^{2}$. Then, the error exponents are as claimed in the theorem, with $E(n)=\frac{1}{6}ct^{2}\beta^{4}$. To conclude, we solve for $\beta$ and substitute the solution $\beta=\sqrt[4]{\frac{6E(n)}{ct^{2}}}$ into Eq. (13). ∎
@@
Given any $\eta>0$ and $10<t<$ 1, and small enough $E(n)>0$. Then for all sufficiently large $n$ there exists a DI code with $E_{1}(n),\,E_{2}(n)\geq E(n)-\frac{3}{n}$ and the following lower bound on the rate in terms of the lower Minkowski dimension:
@@
By definition of the lower Minkowski dimension [Eq. (8)], for any $\eta>0$ there is an $E_{0}>0$ such that for all $E(n)\leq E_{0}$ we have
$\log\left[\Pi_{\sqrt[4]{\frac{6E(n)}{ct^{2}}}}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\right]\geq(\underline{d}_{M}-\eta)\log\sqrt[4]{\frac{ct^{2}}{6E(n)}}.$
The proof is completed by direct substitution of the above expression into Eq. (9) and taking the fourth square root out of the logarithm. ∎
Remark 6.
While the lower bound in Theorem 4 is well suited for small error exponents , it becomes trivial for bigger exponents. Clearly, for $E(n)\geq 2ct^{2}/3$ the packing radius $\beta\geq\sqrt{2}$ is at least the maximum possible distance between any two elements in $\sqrt{\!\widetilde{\mathcal{X}}}$, so the packing number can only be 1. This results on a non-positive (trivial) lower bound on the rate in (9). The reason behind this behaviour and further remarks on the regime of large error exponents can be found in Section IV-D. However, for the main part of the discussion, where the small exponent regime is the relevant one, this will not be a problem.
@@ III-B Upper bound on the rate-reliability function (converse)
For any DI code on block length $n$ with positive error exponents, $E_{1}(n),\,E_{2}(n)\geq E(n)>0$, the rate is upper-bounded
@@
Any DI code satisfies $\frac{1}{2}\|W_{u_{j}}-W_{u_{k}}\|_{1}\geq 1-\lambda_{1}-\lambda_{2}$ for any two different code words $u_{j}\neq u_{k}\in\mathcal{X}^{n}$, thus
$1-\frac{1}{2}\|W_{u_{j}}-W_{u_{k}}\|_{1}\leq 2e^{-E(n)n}.$
Using the relations between fidelity and total variation distance in Eq. (3), this yields
@@
This is the maximum possible fidelity between the output distributions of any two input code words of a good identification code. Our strategy consists in creating a covering such that elements in the same ball have a fidelity higher than the limit stipulated by Eq. (16). In other words, we will ensure that elements in the same ball cannot be well-distinguishable code words, reducing our problem to upper-bounding the cardinality of the covering.
We start with a Euclidean $\frac{\delta}{2}$-covering $\mathcal{X}_{0}$ of the square root output probability set $\sqrt{\!\widetilde{\mathcal{X}}}$ having $|\mathcal{X}_{0}|=\Gamma_{\delta/2}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)=:\Gamma_{\delta/2}$ elements, meaning that for every $x\in\mathcal{X}$ there is a $\xi\in\mathcal{X}_{0}$ such that
$\left|\sqrt{W_{x}}-\sqrt{W_{\xi}}\right|_{2}\leq\frac{\delta}{2}.$
Hence, our input set $\mathcal{X}$ is partitioned according to the $\frac{\delta}{2}$-balls around each $\sqrt{W_{\xi}}$ with $\xi\in\mathcal{X}_{0}$. That is, to each ball corresponds a set of the input sequences that generate only probability distributions in that ball: $\mathcal{X}=\dot{\bigcup}_{\ell=1}^{|\mathcal{X}_{0}|}\mathcal{X}_{\ell}$, such that for all $\ell$ and for any two $x,x^{\prime}\in\mathcal{X}_{\ell}$ we have $|\sqrt{W_{x}}-\sqrt{W_{x^{\prime}}}|_{2}\leq\delta$. We know from Eq. (4) that $1-F(W_{x},W_{x^{\prime}})^{2}\leq|\sqrt{W_{x}}-\sqrt{W_{x^{\prime}}}|_{2}^{2}$. Then, for each $\ell$ and for any two $x,x^{\prime}\in\mathcal{X}_{\ell}$ we have:
$F(W_{x},W_{x^{\prime}})^{2}\geq 1-\delta^{2}.$
In block length $n$, this gives rise to the partition $\mathcal{X}^{n}=\dot{\bigcup}_{\ell^{n}}\mathcal{X}_{\ell^{n}}$, where $\mathcal{X}_{\ell^{n}}:=\mathcal{X}_{\ell_{1}}\times\dots\times\mathcal{X}_{\ell_{n}}$ and $\ell^{n}\in[|\mathcal{X}_{0}|]^{n}$ is the sequence that characterizes the previously defined ball for each letter. This has the property that for all $x^{n},x^{\prime n}\in\mathcal{X}_{\ell^{n}}$,
$F(W_{x^{n}},W_{x^{\prime n}})^{2}\geq(1-\delta^{2})^{n}.$
If we now choose $1-\delta^{2}=e^{-E(n)/2}$, we find that as long as $e^{-E(n)n/2}\geq 4e^{-E(n)n}$ (i.e. $E(n)n/2\geq\ln 4$) the fidelity between elements of the same $\mathcal{X}_{\ell^{n}}$ is bigger than the maximum possible fidelity that the DI code allows. In other words, there cannot be two code words in the same $\mathcal{X}_{\ell^{n}}$. So we can upper bound the number of elements in our code by counting the cardinality of the index $[|\mathcal{X}_{0}|]^{n}$: $N\leq|\mathcal{X}_{0}|^{n}=\left(\Gamma_{\delta/2}\right)^{n}$. Finally, as $\delta=\sqrt{1-e^{-E(n)/2}}$,
$R(n)=\frac{1}{n}\log N\leq\log\left[\Gamma_{\frac{1}{2}{\sqrt{1-e^{-E(n)/2}}}}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\right],$
@@
For any $\eta>0$ and all sufficiently large $n$, for small $E(n)>0$ in the above theorem we can upper-bound the rate of any DI code with error exponents $E_{1}(n),\,E_{2}(n)\geq E(n)$ in terms of the upper Minkowski dimension $\overline{d}_{M}$:
@@
The proof is completed by substituting the above expressions into Eq. (15). ∎

### v4

@@ III Bounding the rate-reliability function
We divide this section into three parts. In III-A, we present a lower bound on the rate-reliability function for DI, expressed in terms of the packing number of a transformed output probability set, where the radius is determined by the error exponents. This is followed by an analysis of the bound in the regime of small error exponents. Similarly, Subsection III-B provides the upper bound using a covering number, along with its analysis for small error exponents. Finally, in III-C we analyse the rate bounds when the error exponents are restricted to certain suitable subsets in the vicinity of 0, raising our lower bound and decreasing our upper bound in the regime of small errors, and illuminating the separate roles of the lower and upper Minkowski dimensions as the objects of merit in pessimistic and optimistic approaches respectively, beyond the capacity value (in the asymptotic limit), for which this separation was first observed in [17].
@@ III-A Lower bound on the rate-reliability function (coding)
We present our lower bound in Theorem 4, and identify the regime where it is more tight in Remark 5. In Corollary 6, we analyse further the regime of small error exponents. Finally, we include an intuitive discussion on the meaning of the bounds. From now on we will use the shorthand $c=1/36K(|\mathcal{Y}|)$.
@@
For any $0<t<1$, and error exponent function $E(n)>0$, there exists a DI code on block length $n$ with error exponents $E_{1}(n)\geq E(n)-\frac{1}{n}$ and $E_{2}(n)\geq E(n)-\frac{3}{n}$, and its rate lower-bounded
$\begin{split}R(n)&\geq(1-t)\log\left[\Pi_{\sqrt[4]{\frac{6E(n)}{ct^{2}}}}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\right]\\ &\phantom{\geq.}-H(t,1-t)-O\left(\frac{\log n}{n}\right),\end{split}$ (17)
with $H(t,1-t)$ the binary entropy of the distribution $\{t,1-t\}$.
@@
Inspired by the code construction in [17], we use the typical sets of input strings as decoding elements, for which we know that Lemmas 2 and 3 give us upper bounds on the errors of first and second kind. We want to impose exponentially decreasing errors, construct a code, and then analyze the resulting rate. Looking at Lemma 2 it is immediate to see that for the error of first kind, the exponential decrease follows by letting $\delta:=\tau\sqrt{n}$, so that $\lambda_{1}\leq 2\exp(-c\tau^{2}n).$
To bound the error of second kind, additional effort is required. We want to use Lemma 3, which requires a minimum separation $1-\frac{1}{2}\|W_{u_{j}}-W_{u_{k}}\|_{1}\leq\epsilon$ between any two different code words $u_{j}\neq u_{k}\in\mathcal{X}^{n}$. That is, the output probability distributions generated by code words need to form a packing with pairwise total variation distance being almost exponentially close to 111. Using the left inequality in Equation (8) first, and the multiplicativity of the fidelity under tensor products after, we have
$1-\frac{1}{2}\|W_{x^{n}}-W_{{x^{\prime}}^{n}}\|_{1}\leq F\left(W_{x^{n}},W_{{x^{\prime}}^{n}}\right)=\prod_{i=1}^{n}F\left(W_{x_{i}},W_{x_{i}^{\prime}}\right).$ (18)
Following [17], we study the negative logarithm of the expression above. Letting $u_{j}=x^{n}=x_{1}\dots x_{n}$ and $u_{k}=x^{\prime n}=x^{\prime}_{1}\dots x^{\prime}_{n}$, we have
@@
Where the second line follows from basic logarithm properties, the third from the leading order of the logarithm when the fidelity is close to 1, and the last line from Equation (9). Notice the important simplification that this achieves: before, we had to construct and analyze a packing of full sequences of length $n$ in total variation distance, and now we just need a letter-wise Euclidean packing to ensure a distance between output probability distributions exponentially close to 1. Let $\mathcal{X}_{0}$ be a maximum-size Euclidean packing in $\sqrt{\!\widetilde{\mathcal{X}}}$ of radius $\beta$ and $\Pi_{\beta}:=\Pi_{\beta}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)=|\mathcal{X}_{0}|$ the corresponding packing number.
Next we choose a maximum-size code $\mathcal{C}_{t}\subset\mathcal{X}_{0}^{n}$ such that the letters are elements of the packing and there is a minimum Hamming distance $d_{H}(x^{n},x^{\prime n})>tn$ between code words $x^{n}\neq x^{\prime n}\in\mathcal{C}_{t}$. Then, in the last sum in Equation (19) at least $tn$ terms are $\geq 4\beta^{2}$ (and could be 0 in the others), therefore
@@
$|\mathcal{C}_{t}|\geq\Pi_{\beta}^{n(1-t)}2^{-nH(t,1-t)}.$ (21)
The Gilbert-Varshamov bound [25, 26] provides a similar (asymptotically equivalent) result, by constructing a linear code over the prime field $\mathbb{F}_{p}$, after reducing $\mathcal{X}_{0}$ to the nearest smaller prime cardinality $p\geq\frac{1}{2}|\mathcal{X}_{0}|$ (Bertrand’s postulate). It only remains to bound the entropy difference $H(W_{x^{n}})-H(W_{x^{\prime n}})$ to apply Lemma 3. As these entropies are in the interval $[0;n\log|\mathcal{Y}|]$, we can just partition the code into $S=\lceil n\log|\mathcal{Y}|\rceil$ parts $\mathcal{C}_{t}^{(s)}$ ($s=1,\ldots,S$) in such a way that all $j\in\mathcal{C}_{t}^{(s)}$ have $H(W_{u_{j}})\in[s-1;s]$, meaning that any two $j,k\in\mathcal{C}_{s}$ satisfy $|H(W_{u_{j}})-H(W_{u_{k}})|\leq 1$. We can now define $\mathcal{C}$ as the largest $\mathcal{C}_{t}^{(s)}$, which by the pigeonhole principle has $|\mathcal{C}|\geq|\mathcal{C}_{t}|/\lceil n\log|\mathcal{Y}|\rceil$ elements. We thus get the rate of the code as
@@
$\begin{split}\lambda_{2}&\leq 2\exp(-c\tau^{2}n)+3\exp(2\tau n-t\beta^{2}n)\\ &\leq 5\exp(-c\tau^{2}n),\end{split}$ (23)
where the last inequality is true if $-c\tau^{2}\geq 2\tau-t\beta^{2}$. Solving the quadratic constraint we obtain
$\tau\leq\frac{-2+\sqrt{4+4ct\beta^{2}}}{2c}=\frac{\sqrt{1+ct\beta^{2}}-1}{c}.$ (24)
We may assume w.l.o.g. $ct\beta^{2}\leq 1$ (see the definition of $c$, the range of $t$ and the meaningful range of $\beta$; otherwise our rate lower bound is trivial), hence by the concavity of the square root the above condition will hold for $\tau=\left(\sqrt{2}-1\right)t\beta^{2}$. Notice that the two error exponents in the present construction are $E_{1}(n)\geq c\tau^{2}-\frac{1}{n}$ and $E_{2}(n)\geq c\tau^{2}-\frac{3}{n}$. Then, they have the form claimed in the theorem if we choose $E(n)=\frac{1}{6}ct^{2}\beta^{4}$. To conclude, we solve for $\beta$ and substitute the solution $\beta=\sqrt[4]{\frac{6E(n)}{ct^{2}}}$ into Equation (22). ∎
Remark 5.
While the lower bound in Theorem 4 is well suited for small error exponents , it becomes trivial for bigger exponents. Clearly, for $E(n)\geq 2ct^{2}/3$ the packing radius $\beta\geq\sqrt{2}$ is at least the maximum possible distance between any two elements in $\sqrt{\!\widetilde{\mathcal{X}}}$, so the packing number can only be 1. This results on a non-positive (trivial) lower bound on the rate in (17). The reason behind this behaviour and further remarks on the regime of large error exponents can be found in Section IV-C. However, for the main part of the discussion, where the small exponent regime is the relevant one as only there we will find super-linearity, this will not be a problem.
@@
[…]
@@
By definition of the lower Minkowski dimension [Equation (16)], for any $\eta>0$ there is an $E_{0}>0$ such that for all $E(n)\leq E_{0}$ we have
$\log\left[\Pi_{\sqrt[4]{\frac{6E(n)}{ct^{2}}}}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\right]\geq(\underline{d}_{M}-\eta)\log\sqrt[4]{\frac{ct^{2}}{6E(n)}}.$ (26)
The proof is completed by direct substitution of the above expression into Equation (17) and taking the fourth square root out of the logarithm. ∎
While the bounds presented in this subsection might look a bit abstract, an important conclusion can be extracted easily. While in transmission the rate-reliability functions converge to the capacity value when the error exponent goes to 0, in deterministic identification it diverges! This behaviour was already hinted in the schematic Figure 1, and it is clear looking at the bound in Corollary 6, where the term $\sim\log(1/E(n))\rightarrow\infty$ when $E(n)\rightarrow 0$. Maybe a bit hidden, we of course have the same effect in Theorem 4, because the packing number of radius $\delta\propto\sqrt[4]{E(n)}\rightarrow 0$ diverges to infinity. So, we have a lower bound that diverges for arbitrary small error exponents, meaning that we can construct a code that achieves infinite linear-scale rate. The correct way to interpret this result is that the rate is not defined in the correct scaling, a faster one is needed. Indeed, we will see in Section IV-A that, in the regime of small error exponents, the rates scaling is linearithmic, and hence, if we define $R(n)=\frac{1}{n\log n}\log N$, the modified reliability function will converge to the (non-trivial) linearithmic capacity lower bound in [17] when $E(n)\rightarrow 0$.
@@ III-B Upper bound on the rate-reliability function (converse)
For any DI code of block length $n$ with positive error exponents, $E_{1}(n),\,E_{2}(n)\geq E(n)>0$, the rate is upper-bounded as follows:
@@
We start by noticing that, given two different codewords $u_{j}\neq u_{k}$ of any DI code, their output distributions have total variation distance bounded. Indeed, we know that there exist a subset $\mathcal{E}_{j}$, the identification test, for which the probability to accept $u_{j}$ is $1-\lambda_{1}$, and $\lambda_{2}$ for any other $u_{k\neq j}$. Therefore, by definition of the total variation distance as a supremum over all possible output subsets [see Equation (7)], it is clear that
$\frac{1}{2}\|W_{u_{j}}-W_{u_{k}}\|_{1}\geq|W_{u_{j}}(\mathcal{E}_{j})-W_{u_{k}}(\mathcal{E}_{j})|\geq 1-\lambda_{1}-\lambda_{2}.$ (28)
Thus,
$1-\frac{1}{2}\|W_{u_{j}}-W_{u_{k}}\|_{1}\leq 2e^{-E(n)n}.$ (29)
Using the relations between fidelity and total variation distance in Equation (8), we get
@@
This is the maximum possible fidelity between the output distributions of any two code words of a good DI code. Now, our strategy consists in creating a covering such that elements in the same ball have a fidelity higher than the limit stipulated by Equation (30). In other words, we will ensure that elements in the same ball cannot be well-distinguishable code words, so that only one code word can be on each ball. Upper-bounding the number of code words (and therefore the rate) is then reduced to upper-bounding the cardinality of the covering.
We start with an Euclidean $\frac{\delta}{2}$-covering $\mathcal{X}_{0}$ of the square root output probability set $\sqrt{\!\widetilde{\mathcal{X}}}$ which has $|\mathcal{X}_{0}|=\Gamma_{\delta/2}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)=:\Gamma_{\delta/2}$ elements, meaning that for every $x\in\mathcal{X}$ there is a $\xi\in\mathcal{X}_{0}$ such that
$\left|\sqrt{W_{x}}-\sqrt{W_{\xi}}\right|_{2}\leq\frac{\delta}{2}.$ (31)
Hence, our input set $\mathcal{X}$ is partitioned according to the $\frac{\delta}{2}$-balls around each $\sqrt{W_{\xi}}$ with $\xi\in\mathcal{X}_{0}$. That is, to each ball $\ell$ corresponds a set of the input sequences $\mathcal{X}_{\ell}$ that generate only probability distributions in that ball. Meaning that for any two $x,x^{\prime}\in\mathcal{X}_{\ell}$ we have $|\sqrt{W_{x}}-\sqrt{W_{x^{\prime}}}|_{2}\leq\delta$. We know from Equation (9) that $1-F(W_{x},W_{x^{\prime}})^{2}\leq|\sqrt{W_{x}}-\sqrt{W_{x^{\prime}}}|_{2}^{2}$. Then,
$F(W_{x},W_{x^{\prime}})^{2}\geq 1-\delta^{2}.$ (32)
The whole input space is partitioned as $\mathcal{X}=\dot{\bigcup}_{\ell=1}^{|\mathcal{X}_{0}|}\mathcal{X}_{\ell}$. In block length $n$, this gives rise to the partition $\mathcal{X}^{n}=\dot{\bigcup}_{\ell^{n}}\mathcal{X}_{\ell^{n}}$, where $\mathcal{X}_{\ell^{n}}:=\mathcal{X}_{\ell_{1}}\times\dots\times\mathcal{X}_{\ell_{n}}$ and $\ell^{n}\in[|\mathcal{X}_{0}|]^{n}$ is the sequence that characterizes the ball for each letter. By virtue of the multiplicativity of the fidelity under tensor products and Equation (32), we have for all $x^{n},x^{\prime n}\in\mathcal{X}_{\ell^{n}}$,
$F(W_{x^{n}},W_{x^{\prime n}})^{2}\geq(1-\delta^{2})^{n}.$ (33)
If we now choose $1-\delta^{2}=e^{-E(n)/2}$, we find that as long as $e^{-E(n)n/2}\geq 4e^{-E(n)n}$ (i.e. $E(n)n/2\geq\ln 4$) the fidelity between elements of the same $\mathcal{X}_{\ell^{n}}$ is bigger than the maximum possible fidelity that the DI code allows. In other words, there cannot be two code words in the same $\mathcal{X}_{\ell^{n}}$. So we can upper bound the number of elements in our code by counting the cardinality of the index $[|\mathcal{X}_{0}|]^{n}$: $N\leq|\mathcal{X}_{0}|^{n}=\left(\Gamma_{\delta/2}\right)^{n}$ with the chosen value $\delta=\sqrt{1-e^{-E(n)/2}}$. We get
$R(n)=\frac{1}{n}\log N\leq\log\left[\Gamma_{\frac{1}{2}{\sqrt{1-e^{-E(n)/2}}}}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\right],$ (34)
@@
Let $\eta>0$ and a small enough $E(n)>0$ be given. Then, for all sufficiently large $n$, the rate of any DI code with error exponents $E_{1}(n),\,E_{2}(n)\geq E(n)$ can be upper-bounded in terms of the upper Minkowski dimension $\overline{d}_{M}$ as
@@
The proof is completed by substituting the above expressions into Equation (27). ∎